# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ Appendix A Kadanoff-Baym Formalism
$=-t_{1^{\prime}3^{\prime}}c_{1^{\prime}}(\tau)\delta_{3^{\prime},5}\delta_{\tau,\tau_{5}}+\frac{1}{2}U_{1^{\prime}2^{\prime}3^{\prime}4^{\prime}}\left(c_{3^{\prime}}^{\dagger}(\tau)\delta_{5,4^{\prime}}-\delta_{5,3^{\prime}}c_{4^{\prime}}^{\dagger}(\tau)\right)\delta_{\tau,\tau_{5}}c_{2^{\prime}}(\tau)c_{1^{\prime}}(\tau)$ (77)
$=-t_{1^{\prime}3^{\prime}}c_{1^{\prime}}(\tau_{5})\delta_{3^{\prime},5}+U_{1^{\prime}2^{\prime}3^{\prime}4^{\prime}}c_{3^{\prime}}^{\dagger}(\tau_{5})\delta_{4^{\prime},5}c_{2^{\prime}}(\tau_{5})c_{1^{\prime}}(\tau_{5})\,,$ (78)
where we exploited the crossing symmetry of $U$ and relabeled the summation indices in the last line. From this we find the equation of motion for the Green’s function as
@@
$\partial_{\tau_{a}}G_{a,b,\phi}(\tau_{a},\tau_{b})=\phi_{a1^{\prime}}G_{1^{\prime}b}(\tau_{1}^{\prime},\tau_{b})+\delta_{a,b}-t_{1a}\braket{c_{1}(\tau_{a})c^{\dagger}_{b}(\tau_{b})}+U_{1^{\prime}2^{\prime}3^{\prime}a}\braket{c_{3^{\prime}}^{\dagger}(\tau_{a})c_{2^{\prime}}(\tau)c_{1^{\prime}}(\tau_{a})c^{\dagger}_{b}(\tau_{b})}\,.$ (79)
By comparing with the Schwinger Dyson equation we extract
$\Sigma_{1,1^{\prime}}(\tau_{1},\tau_{1}^{\prime})G_{1^{\prime},3}(\tau_{1}^{\prime},\tau_{3})=U_{4^{\prime},2^{\prime},5^{\prime},1}\braket{c_{5^{\prime}}^{\dagger}(\tau_{1})c_{2^{\prime}}(\tau_{1})c_{4^{\prime}}(\tau_{1})c^{\dagger}_{3}(\tau_{3})}\,,$ (80)
which is the central equation for all following derivations. Note that we could have analogously taken the initial derivative w.r.t $\tau_{b}$ which would result in
$\Sigma_{1,1^{\prime}}(\tau_{1},\tau_{1}^{\prime})G_{1^{\prime},3}(\tau_{1}^{\prime},\tau_{3})=U_{4^{\prime},2^{\prime},5^{\prime},1}\braket{c_{5^{\prime}}^{\dagger}(\tau_{3})c_{2^{\prime}}(\tau_{3})c_{4^{\prime}}(\tau_{1})c^{\dagger}_{3}(\tau_{3})}\,.$ (81)
@@ A.1 Self-energy
We can re-express the implicit equation for the self-energy as an explicit one, the Schwinger-Dyson equation:
$\Sigma_{1,1^{\prime}}(\tau_{1},\tau_{1}^{\prime})G_{1^{\prime},3}(\tau_{1}^{\prime},\tau_{3})$ $=U_{4^{\prime},2^{\prime},5^{\prime},1}\braket{c_{5^{\prime}}^{\dagger}(\tau_{1})c_{2^{\prime}}(\tau_{1})c_{4^{\prime}}(\tau_{1})c^{\dagger}_{3}(\tau_{3})}$ (82)
$=U_{4^{\prime},2^{\prime},5^{\prime},1}\left(\chi_{2^{\prime},4^{\prime},5^{\prime},3}(\tau_{1},\tau_{1},\tau_{1},\tau_{3})-G_{2^{\prime},5^{\prime}}(\tau_{1},\tau_{1})G_{4^{\prime},3}(\tau_{1},\tau_{3})\right)$ (83)
$=U_{4^{\prime},2^{\prime},5^{\prime},1}\Big{(}\chi_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma_{b^{\prime},a^{\prime},c^{\prime},d^{\prime}}\ G_{c^{\prime},3}(\tau_{5}\tau_{3})G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})$ (84)
$\quad+G_{2^{\prime},3}(\tau_{1},\tau_{3})G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})-G_{2^{\prime},5^{\prime}}(\tau_{1},\tau_{1})G_{4^{\prime},3}(\tau_{1},\tau_{3})\Big{)}$ (85)
$\Leftrightarrow\Sigma_{1,5}(\tau_{1},\tau_{5})$ $=U_{4^{\prime},2^{\prime},5^{\prime},1}\Big{(}\chi_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma_{b^{\prime}a^{\prime}5d^{\prime}}\ G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})$ (86)
$\quad+2G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{2^{\prime},5}\delta_{\tau_{1},\tau_{5}}\Big{)}\,.$ (87)
Here, the latter term corresponds to the Hartree-Fock selfenergy, while the former is the Schwinger-Dyson equation. At this point, we are working in diagrammatic spin-space. However, the susceptibility and the vertex are only known in the physical spin-space. Furthermore, we can utilize that the self-energy and Green’s function are spin diagonal. Hence, we pull the spin-dependence out of the combined index
$\Sigma_{1,5}(\tau_{1},\tau_{5})\delta_{s_{1},\sigma_{5}}$ $=(U_{4^{\prime},2^{\prime},5^{\prime},1}\delta_{s_{4}^{\prime},\sigma_{5}^{\prime}}\delta_{s_{2}^{\prime},s_{1}}-U_{4^{\prime}2^{\prime}15^{\prime}}\delta_{s_{4}^{\prime},s_{1}}\delta_{s_{2}^{\prime},\sigma_{5}^{\prime}})$ (88)
$\quad\Big{(}\frac{1}{4}S^{i^{\prime}}_{s_{2}^{\prime},\sigma_{5}^{\prime}}S^{i^{\prime}}_{\sigma_{a}^{\prime},\sigma_{b}^{\prime}}\chi^{S^{i^{\prime}}S^{i^{\prime}}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})S^{j^{\prime}}_{\sigma_{b}^{\prime},\sigma_{a}^{\prime}}S^{j^{\prime}}_{\sigma_{5},\sigma_{d}^{\prime}}\Gamma^{S^{j^{\prime}}S^{j^{\prime}}}_{b^{\prime}a^{\prime}5d^{\prime}}G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\delta_{s_{4}^{\prime},\sigma_{d}^{\prime}}$
$\quad+\delta_{2^{\prime},5}\delta_{\tau_{1},\tau_{5}}G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{s_{4}^{\prime},\sigma_{5}^{\prime}}-G_{2^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{4^{\prime},5}\delta_{\tau_{1},\tau_{5}}\delta_{s_{2}^{\prime},\sigma_{5}^{\prime}}\Big{)}$
$=2U_{4^{\prime},2^{\prime},5^{\prime},1}G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{2^{\prime},5}\delta_{\tau_{1},\tau_{5}}-U_{4^{\prime}2^{\prime}15^{\prime}}G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{2^{\prime},5}\delta_{\tau_{1},\tau_{5}}$ (89)
$\quad+\frac{1}{4}U_{4^{\prime},2^{\prime},5^{\prime},1}G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\chi^{S^{i^{\prime}}S^{i^{\prime}}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma^{S^{j^{\prime}}S^{j^{\prime}}}_{b^{\prime}a^{\prime}5d^{\prime}}S^{i^{\prime}}_{s_{1},\sigma_{d}^{\prime}}S^{i^{\prime}}_{\sigma_{a}^{\prime},\sigma_{b}^{\prime}}S^{j^{\prime}}_{\sigma_{b}^{\prime},\sigma_{a}^{\prime}}S^{j^{\prime}}_{\sigma_{5},\sigma_{d}^{\prime}}$
$\quad-\frac{1}{4}U_{4^{\prime}2^{\prime}15^{\prime}}G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\chi^{S^{i^{\prime}}S^{i^{\prime}}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma^{S^{j^{\prime}}S^{j^{\prime}}}_{b^{\prime}a^{\prime}5d^{\prime}}S^{i^{\prime}}_{\sigma_{5}^{\prime},\sigma_{5}^{\prime}}\delta_{i^{\prime},j^{\prime}}2S^{j^{\prime}}_{\sigma_{5},s_{1}}$
$=2U_{4^{\prime},2^{\prime},5^{\prime},1}G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{2^{\prime},5}\delta_{\tau_{1},\tau_{5}}-U_{4^{\prime}2^{\prime}15^{\prime}}G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{2^{\prime},5}\delta_{\tau_{1},\tau_{5}}$ (90)
$\quad+U_{4^{\prime},2^{\prime},5^{\prime},1}G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\chi^{S^{i^{\prime}}S^{i^{\prime}}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma^{S^{i^{\prime}}S^{i^{\prime}}}_{b^{\prime}a^{\prime}5d^{\prime}}$
$\quad-U_{4^{\prime}2^{\prime}15^{\prime}}G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\chi^{S^{0}S^{0}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma^{S^{0}S^{0}}_{b^{\prime}a^{\prime}5d^{\prime}}$
$=(2U_{4^{\prime},5,5^{\prime},1}-U_{4^{\prime}515^{\prime}})G_{4^{\prime},5^{\prime}}(\tau_{1},\tau_{1})\delta_{\tau_{1},\tau_{5}}$ (91)
$\quad+3U_{4^{\prime},2^{\prime},5^{\prime},1}G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\chi^{S^{z}S^{z}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma^{S^{z}S^{z}}_{b^{\prime}a^{\prime}5d^{\prime}}$
$\quad+(2U_{4^{\prime}2^{\prime}15^{\prime}}-U_{4^{\prime},2^{\prime},5^{\prime},1})G_{4^{\prime},d^{\prime}}(\tau_{1},\tau_{5})\chi^{S^{0}S^{0}}_{2^{\prime},a^{\prime},5^{\prime},b^{\prime}}(\tau_{1},\tau_{5},\tau_{1},\tau_{5})\Gamma^{S^{0}S^{0}}_{b^{\prime}a^{\prime}5d^{\prime}}\,.$
There are two internal consistency checks - one for each flavor of TPSC, which is the usual internal consistency check between single and two-particle objects in the pairwise equal position, time and orbital limit:
$\Sigma_{1,1^{\prime}}G_{1^{\prime},1}$ $=\Big{(}D_{o_{4},o_{1}}\sum_{s_{4}}\bar{\delta}_{s_{1},s_{4}}^{o_{1},o_{4}}\braket{c_{o_{4},s_{4}}^{\dagger}c_{o_{1},s_{1}}c_{o_{4},s_{4}}c_{1}^{\dagger}}+C_{o_{1},o_{4}}\sum_{s_{4}}\bar{\delta}_{s_{1},s_{4}}^{o_{1},o_{4}}\braket{c_{o_{4},s_{4}}^{\dagger}c_{o_{4},s_{1}}c_{o_{1},s_{4}}c_{1}^{\dagger}}$ (92)
$\quad+\bar{\delta}_{s_{1},s_{4}}P_{o_{2},o_{1}}\sum_{s_{4}}\braket{c_{o_{1},s_{4}}^{\dagger}c_{o_{2},s_{1}}c_{o_{2},s_{4}}c_{1}^{\dagger}}\Big{)}.$
Both TPSC3 and TPSC5 fulfill this up to numerical precision, as long as the self consistency loop for the spin channel TPEV’s and the subsequent TPEV fitting in the charge channel have a solution which is physical. A second consistency check is the equivalence between TPSC3 and TPSC5.

### v3

@@ Appendix A Kadanoff-Baym Formalism
$\quad-\frac{1}{2}U_{o_{1^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{4^{\prime}}}^{s_{1^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{4^{\prime}}}\left(\left(c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}^{\prime})\delta_{o_{4^{\prime}},o_{5}}\delta_{s_{4^{\prime}},s_{5}}\delta_{\boldsymbol{r}^{\prime},\boldsymbol{r}_{5}}-\delta_{o_{3^{\prime}},o_{5}}\delta_{s_{3^{\prime}},s_{5}}\delta_{\boldsymbol{r}^{\prime},\boldsymbol{r}_{5}}c^{\dagger}_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{r}^{\prime})\right)\right.$
$\qquad\left.c_{o_{2^{\prime}},s_{2^{\prime}}}(\boldsymbol{r}^{\prime})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}^{\prime})\right)_{H}$
$=-t_{o_{1^{\prime}},o_{5}}^{s_{1^{\prime}},s_{5}}(\boldsymbol{r}_{1}^{\prime},\boldsymbol{r}_{5})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}_{1}^{\prime},\tau_{5})$ (69)
$\quad+\underbrace{\frac{1}{2}\left(U_{o_{1^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{5}}^{s_{1^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{5}}-U_{o_{1^{\prime}},o_{2^{\prime}},o_{5},o_{3^{\prime}}}^{s_{1^{\prime}},s_{2^{\prime}},s_{5},s_{3^{\prime}}}\right)}_{=\Gamma_{o_{1^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{5}}^{s_{1^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{5}}}c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{r}_{5},\tau_{5})c_{o_{2^{\prime}}s_{2^{\prime}}}(\boldsymbol{r}_{5},\tau_{5})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{r}_{5},\tau_{5})\;,$
where we used that the Hamiltonian commutes with itself at any time to pull the time evolution operator out of the commutator. The time evolution operators are indicated by the subscript $H$ on the respective outermost commutator or bracket. Furthermore, we remark that in the case of an anti-symmetrized interaction $U$ we have $\Gamma=\alpha U$, where the prefactor $\alpha$ depends on the exact definition of the anti-symmetrization.
In the next step, we derive the equation of motion of the Green’s function, where we first have to perform a partial time ordering such that the derivative commutes with the time ordering operator (for brevity, we will use numbers as a shorthand for a collection of all quantum numbers, whenever they all do share the same index, i.e. $1\equiv(o_{1},s_{1},\boldsymbol{r}_{1},\tau_{1})$),
$-Z_{\phi}\partial_{\tau_{1}}G_{1,3}$ $=\partial_{\tau_{1}}\left(\braket{T_{\tau}S[\phi]c_{1}c^{\dagger}_{3}}\Theta(\tau_{1}-\tau_{3})-\braket{T_{\tau}S[\phi]c^{\dagger}_{3}c_{1}}\Theta(\tau_{3}-\tau_{1})\right)$ (70)
$=\braket{\partial_{\tau_{1}}T_{\tau}S[\phi]c_{1}c^{\dagger}_{3}}\Theta(\tau_{1}-\tau_{3})-\braket{\partial_{\tau_{1}}T_{\tau}S[\phi]c^{\dagger}_{3}c_{1}}\Theta(\tau_{3}-\tau_{1})$ (71)
$\quad+\underbrace{\braket{T_{\tau}(c_{1}c^{\dagger}_{3}+c^{\dagger}_{3}c_{1})S[\phi]}\delta(\tau_{1}-\tau_{3})}_{=\delta_{1,3}Z[\phi]}.$
To evaluate the time derivative we first have to time order the exponential and creation/annihilation operators partially, leaving out the time ordering w.r.t the argument of the annihilation operator of the exponential. We perform this fist for the first contribution
[…]
$=\Braket{T_{\tau}S[\phi]\left(\partial_{\tau_{1}}c_{1}\right)c^{\dagger}_{3}}-\Braket{T_{\tau}S[\phi]\phi_{1^{\prime},1}c_{1^{\prime}}c^{\dagger}_{3}}\,.$ (74)
The second contribution follows analogously and both of them can be recombined by utilizing the time ordering. From this we directly find
@@
$\partial_{\tau_{1}}G_{1,3}$ $=-\frac{1}{Z_{\phi}}\braket{T_{\tau}S[\phi]\left(\partial_{\tau_{1}}c_{1}\right)c^{\dagger}_{3}}-\phi_{1^{\prime},1}G_{1^{\prime},3}-\delta_{1,3}.$ (75)
Now, let us insert the equation of motion for the annihilation operator,
$\partial_{\tau_{1}}G_{1,3}$ $=-\Gamma_{o_{1^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{1^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{2^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{1^{\prime}},s_{1^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}_{\phi}$ (76)
$-t_{1^{\prime},1}\delta_{\tau_{1}^{\prime},\tau_{1}}G_{1^{\prime},3}-\phi_{1^{\prime},1}G_{1^{\prime},3}-\delta_{1,3}.$ (77)
By setting the interaction and the sources to zero, we identify the non-interacting Green’s function as
$(G^{(0)})^{-1}_{1,3}=(-\partial_{\tau_{1}}\delta_{1,3}-t_{1,3}\delta_{\tau_{1},\tau_{3}}).$ (78)
Furthermore, we identify the self-energy as
$\Sigma_{o_{1},o_{1^{\prime}}}^{s_{1},s_{1^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1}^{\prime})$ $G_{o_{1^{\prime}},o_{3}}^{s_{1^{\prime}},s_{3}}(\boldsymbol{\tau}_{1}^{\prime},\boldsymbol{\tau}_{3})$ (79)
$=\Gamma_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{2^{\prime}}s_{2^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{3},s_{3}}^{\dagger}(\boldsymbol{\tau}_{3})}_{\phi},$
such that we arrive at the following Dyson equation for the interacting Green’s function:
$G^{-1}_{1,3}=(G^{(0)})^{-1}_{1,3}-\phi_{3,1}-\Sigma_{1,3}\,.$ (80)
Note that due to the definition of the source-field term, there is an index swap in this equation compared to all other quantities appearing on the right hand side of the equation.
The two-particle expectation values on the RHS of Eq. (79) are directly linked to the generalized two-particle susceptibility, which is defined as the second functional derivative of $\mathcal{G}[\phi]$:
$\chi_{o_{1}..o_{4}}^{s_{1}..s_{4}}(\boldsymbol{\tau}_{1}..\boldsymbol{\tau}_{4})$ $=\frac{\delta G^{s_{1},s_{3}}_{o_{1},o_{3};\phi}(\boldsymbol{\tau}_{1}\boldsymbol{\tau}_{3})}{\delta\phi^{s_{2},s_{4}}_{o_{2},o_{4};\phi}(\boldsymbol{\tau}_{2}\boldsymbol{\tau}_{4})}$ (81)
$=\braket{T_{\tau}c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{\tau}_{3})c^{\dagger}_{o_{4},s_{4}}(\boldsymbol{\tau}_{4})c_{o_{2},s_{2}}(\boldsymbol{\tau}_{2})c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})}_{\phi}$
$\quad-\braket{T_{\tau}c_{o_{1},s_{1}}(\boldsymbol{\tau}_{1})c^{\dagger}_{o_{3},s_{3}}(\boldsymbol{\tau}_{3})}_{\phi}\braket{T_{\tau}c_{o_{2},s_{2}}(\boldsymbol{\tau}_{2})c^{\dagger}_{o_{4},s_{4}}(\boldsymbol{\tau}_{4})}_{\phi}\,,$
while the two-particle vertex is given by
$\Gamma_{o_{1}..o_{4}}^{s_{1}..s_{4}}(\boldsymbol{\tau}_{1}..\boldsymbol{\tau}_{4})=\frac{\delta\Sigma_{o_{1},s_{4}}^{s_{1},s_{4}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{4})}{\delta G_{o_{3},s_{2}}^{s_{3},s_{2}}(\boldsymbol{\tau}_{3},\boldsymbol{\tau}_{2})}\,.$ (82)
With these definitions we obtain the Bethe-Salpeter equation connecting the interacting and non-interacting susceptibilities:
00 0 $=\frac{\delta(G^{-1}_{1,3^{\prime}}G_{3^{\prime},3})}{\delta{\phi_{2,4}}}$ (83)
$=G^{-1}_{1,3^{\prime}}\frac{\delta(G_{3^{\prime},3})}{\delta{\phi_{2,4}}}+\frac{\delta(G^{-1;0}_{1,3^{\prime}}-\phi_{3^{\prime},1}-\Sigma_{1,3^{\prime}})}{\delta{\phi_{2,4}}}G_{3^{\prime},3}$ (84)
$\Leftrightarrow\chi_{1,2,3,4}$ $=G_{1,1^{\prime}}\frac{\delta(\phi_{3^{\prime},1^{\prime}}+\Sigma_{1^{\prime},3^{\prime}})}{\delta{\phi_{2,4}}}G_{3^{\prime},3}$ (85)
$=G_{1,4}G_{2,3}+G_{1,1^{\prime}}\frac{\delta\Sigma_{1^{\prime},3^{\prime}}}{\delta{\phi_{2,4}}}G_{3^{\prime},3}$ (86)
$=G_{1,4}G_{2,3}+G_{1,1^{\prime}}\frac{\delta\Sigma_{1^{\prime},3^{\prime}}}{\delta{G_{2^{\prime},4^{\prime}}}}\frac{\delta G_{2^{\prime},4^{\prime}}}{\delta{\phi_{2,4}}}G_{3^{\prime},3}$ (87)
$=G_{1,4}G_{2,3}+G_{1,1^{\prime}}G_{3^{\prime},3}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}\chi_{2^{\prime},2,4^{\prime},4}$ (88)
$=\chi^{0}_{1,2,3,4}+\chi^{0}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}\chi_{2^{\prime},2,4^{\prime},4}$ (89)
$=\chi^{0}_{1,2,3,4}+\chi_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}\chi^{0}_{2^{\prime},2,4^{\prime},4}\,.$ (90)
For numerical purposes it is often beneficial to reorder the indices such that the equations can be written as a matrix-matrix product.
In an $SU(2)$ symmetric system, we can furthermore use that all occurring quantities are diagonal in physical spin-space, thus resulting in spin-channel-specific equations which read (pulling out spin indices from the numbers)
$\chi^{S^{i^{\prime}},S^{j^{\prime}}}_{1,2,3,4}$ $=(\chi^{0}_{1,2,3,4}+\chi^{0}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}\chi_{2^{\prime},2,4^{\prime},4})\sigma^{i}_{s_{1},s_{3}}\sigma^{j}_{s_{2},s_{4}}$ (91)
$\Leftrightarrow\chi_{1,2,3,4}^{S^{z},S^{z}}$ $=\chi_{1,2,3,4}^{0;S^{z},S^{z}}+\chi^{0;S^{z},S^{z}}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}^{S^{z},S^{z}}\chi_{2^{\prime},2,4^{\prime},4}^{S^{z},S^{z}}$ (92)
$\Leftrightarrow\chi_{1,2,3,4}^{S^{0},S^{0}}$ $=\chi_{1,2,3,4}^{0;S^{0},S^{0}}-\chi^{0;S^{0},S^{0}}_{1,3^{\prime},3,1^{\prime}}\Gamma_{1^{\prime},3^{\prime},4^{\prime},2^{\prime}}^{S^{0},S^{0}}\chi_{2^{\prime},2,4^{\prime},4}^{S^{0},S^{0}}\,,$ (93)
where $S_{x}S_{x}$, $S_{y}S_{y}$ and $S_{z}S_{z}$ are related by $SU(2)$ symmetry and $S^{0}$ indicates the case in which $\sigma^{i}=\sigma^{0}_{s_{1},s_{2}}=\delta_{s_{1},s_{2}}$.
@@ A.1 Self-energy
In the following we will derive the equations required for the level-1 self-energy evaluation. For the following calculations, we assume an anti-symmetrized interaction tensor
$\Sigma_{o_{1},o_{3}}^{s_{1},s_{3}}$ $(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3})=U_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}\braket{T_{\tau}c^{\dagger}_{o_{3^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{5}^{\prime},s_{5}^{\prime}}^{\dagger}(\boldsymbol{\tau}_{5}^{\prime})c_{o_{2^{\prime}}s_{2^{\prime}}}(\boldsymbol{\tau}_{1})c_{o_{4^{\prime}},s_{4^{\prime}}}(\boldsymbol{\tau}_{1})}_{\phi}$ (94)
×$\qquad\qquad\qquad\times(G^{-1})_{o_{5}^{\prime},o_{3}}^{s_{5}^{\prime},s_{3}}(\boldsymbol{\tau}_{5}^{\prime},\boldsymbol{\tau}_{3})$
$=U_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}\left(\chi_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{5^{\prime}}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{5^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{5^{\prime}})\right.$ (95)
$\left.\qquad\qquad\qquad-G_{o_{4^{\prime}},o_{3^{\prime}}}^{s_{4^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1})G_{o_{2^{\prime}},o_{5^{\prime}}}^{s_{2^{\prime}},s_{5^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{5}^{\prime})\right)(G^{-1})_{o_{5}^{\prime},o_{3}}^{s_{5}^{\prime},s_{3}}(\boldsymbol{\tau}_{5}^{\prime},\boldsymbol{\tau}_{3})$
$=U_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}\Big{(}\chi_{o_{4^{\prime}},o_{c^{\prime}},o_{3^{\prime}},o_{a^{\prime}}}^{s_{4^{\prime}},s_{c^{\prime}},s_{3^{\prime}},s_{a^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{c^{\prime}},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{a^{\prime}})$ (96)
$\quad\quad\quad\quad\quad\Gamma_{o_{a^{\prime}},o_{c^{\prime}},o_{d^{\prime}},o_{b^{\prime}}}^{s_{a^{\prime}},s_{c^{\prime}},s_{d^{\prime}},s_{b^{\prime}}}(\boldsymbol{\tau}_{a^{\prime}},\boldsymbol{\tau}_{c^{\prime}},\boldsymbol{\tau}_{d^{\prime}},\boldsymbol{\tau}_{b^{\prime}})G_{o_{2^{\prime}},o_{d^{\prime}}}^{s_{2^{\prime}},s_{d^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{d^{\prime}})\delta_{o_{b^{\prime}},o_{3}}\delta_{s_{b^{\prime}},s_{3}}\delta_{\boldsymbol{\tau}_{b^{\prime}},\boldsymbol{\tau}_{3}}$
$\quad+\delta_{o_{4^{\prime}},o_{3}}\delta_{s_{4^{\prime}},s_{3}}\delta_{\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3}}G_{o_{2^{\prime}},o_{3^{\prime}}}^{s_{2^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1})-G_{o_{4^{\prime}},o_{3^{\prime}}}^{s_{4^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1})\delta_{o_{2^{\prime}},o_{3}}\delta_{s_{2^{\prime}},s_{3}}\delta_{\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3}}\Big{)}$
$=U_{o_{4^{\prime}},o_{2^{\prime}},o_{3^{\prime}},o_{1}}^{s_{4^{\prime}},s_{2^{\prime}},s_{3^{\prime}},s_{1}}$ (97)
$\quad\quad\Big{(}\chi_{o_{4^{\prime}},o_{c^{\prime}},o_{3^{\prime}},o_{a^{\prime}}}^{s_{4^{\prime}},s_{c^{\prime}},s_{3^{\prime}},s_{a^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{c^{\prime}},\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{a^{\prime}})\Gamma_{o_{a^{\prime}},o_{c^{\prime}},o_{d^{\prime}},o_{3}}^{s_{a^{\prime}},s_{c^{\prime}},s_{d^{\prime}},s_{3}}(\boldsymbol{\tau}_{a^{\prime}},\boldsymbol{\tau}_{c^{\prime}},\boldsymbol{\tau}_{d^{\prime}},\boldsymbol{\tau}_{3})G_{o_{2^{\prime}},o_{d^{\prime}}}^{s_{2^{\prime}},s_{d^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{d^{\prime}})$
$\quad\quad\quad\quad+2\delta_{o_{4^{\prime}},o_{3}}\delta_{s_{4^{\prime}},s_{3}}\delta_{\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{3}}G_{o_{2^{\prime}},o_{3^{\prime}}}^{s_{2^{\prime}},s_{3^{\prime}}}(\boldsymbol{\tau}_{1},\boldsymbol{\tau}_{1})\Big{)}\,.$
Here, the latter term corresponds to the Hartree-Fock selfenergy, while the former corresponds to the Schwinger-Dyson equation.